# Two questions on Kneser colorings

### v1

@@ Abstract
In this paper, we investigate two questions on Kneser graphs $KG_{n,k}$. First, we prove that the union of $s$ non-trivial intersecting families in (${[n]\choose k}$ has size at most (${n\choose k}-{n-s\choose k}$ for all sufficiently large $n$ that satisfy $n>(2+\epsilon)k^{2}$ with $\epsilon>0$. We provide an example that shows that this result is essentially tight for the number of colors close to $\chi(KG_{n,k})=n-2k+2$. We also improve the result of Bulankina and Kupavskii on the choice chromatic number, showing that it is at least $\frac{1}{16}n\log n$ for all $k<\sqrt{n}$ and $n$ sufficiently large.
@@ 1 Introduction
The following generalization of the Erdős–Ko–Rado theorem was proposed by Erdős [3]: for given $n,k,s$, determine, what is the largest size of a family ($\mathcal{F}\subset{[n]\choose k}$ such that $\mathcal{F}$ is a union of at most $s$ intersecting families. It is a natural conjecture to make that for triples of parameters $(n,k,s)$ where $n$ is somewhat large in comparison to $k,s$, the extremal example is a union of $s$ stars. Erdős indeed conjectured this for all $n\geq 2k+t-1$. In this case, ||=($|\mathcal{F}|={n\choose k}-{n-s\choose k}$. Frankl and Fűredi [5] showed that the conjecture is true for $s=2$ and $n>ck$ with $c\approx 2.62$. However, it is not true in general. This was shown for $n=2k+t-1$ by Hilton and Füredi. Frankl and Füredi also gave an example which shows that the extreme is different even for $s=2$, when $n\leq 2k+c_{0}\sqrt{k}$ for $n\geq n_{0}$ and some positive constant $c_{0}$. Ellis and Lifshitz [2] recently proved that for constant $s$ and $n>2k+C(s)k^{2/3}$ the union of stars is optimal, where $C(s)$ is a function of $s$.
@@
The proofs of Theorems 1 and 3 both rely on structural properties of intersecting families that were studied in the paper of Kiselev and the second author [12].
@@ 2.1 Preliminaries
We call families $\mathcal{A},\mathcal{B}\subset 2^{X}$ cross-intersecting, if $A\cap B\neq\emptyset$ for any $A\in\mathcal{A},B\in\mathcal{B}$. We say that $C\subset X$ covers family $\mathcal{F}\subset 2^{X}$ if $C\cap A\neq\emptyset$ for any $A\in\mathcal{F}$. We denote the size of the smallest cover of $\mathcal{F}$ by $\tau(\mathcal{F})$. We say that a family $\mathcal{H}$ set-covers a family $\mathcal{F}$ if for each $F\in\mathcal{F}$ there exists $H\in\mathcal{H}$ such that $H\subset F$. For a graph $G$ denote $\mathcal{I}_{k}(G)$ the family of independent $k$-element sets of vertices of $G$. In the proofs of Theorems 1 and 3, we need the following result from [12]:
@@
We construct $\mathcal{H}$ inductively. Put $\mathcal{H}_{0}=\emptyset$. For each $l\in[0,\sqrt{k}-1]$, we construct a set-cover $\mathcal{H}_{l+1}\subset{[n]\choose l+1}$ for $\mathcal{F}$ from the set-cover $\mathcal{H}_{l}\subset{[n]\choose l}$ that we assume to have constructed. We also assume that $|\mathcal{H}_{l}|\leq k^{l}$. Recall that $\mathcal{H}_{l}$ being a set-cover means that $\mathcal{F}\subset\bigcup_{H\in\mathcal{H}_{l}}\mathcal{F}(H)$. Thus, it should be clear that $\mathcal{H}_{0}$ satisfies both the condition on size and on being a set-cover. For each $H\in\mathcal{H}_{l}$ consider the set $G_{H}\in\mathcal{G}^{\times}$ disjoint with $H$. Such set exists since $|H|=l<\sqrt{k}$ and $\tau(\mathcal{G}^{\times})\geq\sqrt{k}$. Put $\mathcal{H}_{l+1}=\{H\cup x:H\in\mathcal{H}_{l},x\in G_{H}\}$. Then it is easy to see that $|\mathcal{H}_{l+1}|\leq t|\mathcal{H}_{l}|\leq k^{l+1}$ and
$\mathcal{F}\subset\bigcup_{H\in\mathcal{H}_{l+1}}\mathcal{F}(G).$
@@ 2.2 Proof of Theorem 1
    $\mathcal{N}_{k}(\bar{G})\geq\frac{2(k-1)\gamma-(k-2)}{k}|V(\bar{G})|\mathcal{N}_{k-1}(\bar{G}),$
@@
    $\mathcal{N}_{3}(\bar{G})\geq\frac{2\cdot 2\gamma-1}{3}|V(\bar{G})|\mathcal{N}_{2}(\bar{G}).$

### v2

@@ Abstract
In this paper, we investigate two questions on Kneser graphs $KG_{n,k}$. First, we prove that the union of $s$ intersecting families in (${[n]\choose k}$ has size at most (${n\choose k}-{n-s\choose k}$ for all sufficiently large $n$ that satisfy $n>(2+\epsilon)k^{2}+s$ with $\epsilon>0$. We provide an example that shows that this result is essentially tight for the number of colors close to $\chi(KG_{n,k})=n-2k+2$. We also improve the result of Bulankina and Kupavskii on the choice chromatic number, showing that it is at least $\frac{1}{25}n\log n$ for all $k<\sqrt{n}$ and $n$ sufficiently large.
@@ 1 Introduction
The following generalization of the Erdős–Ko–Rado theorem was proposed by Erdős [3]: for given $n,k,s$, determine, what is the largest size of a family ($\mathcal{F}\subset{[n]\choose k}$ such that $\mathcal{F}$ is a union of at most $s$ intersecting families. It is a natural conjecture to make that for triples of parameters $(n,k,s)$ where $n$ is somewhat large in comparison to $k,s$, the extremal example is a union of $s$ stars. Erdős indeed conjectured this for all $n\geq 2k+s-1$. In this case, ||=($|\mathcal{F}|={n\choose k}-{n-s\choose k}$. Frankl and Füredi [5] showed that the conjecture is true for $s=2$ and $n>ck$ with $c\approx 2.62$. However, it is not true in general. This was shown for $n=2k+s-1$ by Hilton and Füredi. Frankl and Füredi also gave an example which shows that the extreme is different even for $s=2$, when $n\leq 2k+c_{0}\sqrt{k}$ for $n\geq n_{0}$ and some positive constant $c_{0}$. Ellis and Lifshitz [2] recently proved that for constant $s$ and $n>2k+C(s)k^{2/3}$ the union of stars is optimal, where $C(s)$ is a function of $s$.
@@
The proofs of Theorems 1 and 4 both rely on structural properties of intersecting families that were studied in the paper of Kiselev and the second author [12].
@@ 2.1 Preliminaries
We call families $\mathcal{A},\mathcal{B}\subset 2^{X}$ cross-intersecting, if $A\cap B\neq\emptyset$ for any $A\in\mathcal{A},B\in\mathcal{B}$. We say that $C\subset X$ covers family $\mathcal{F}\subset 2^{X}$ if $C\cap A\neq\emptyset$ for any $A\in\mathcal{F}$. We denote the size of the smallest cover of $\mathcal{F}$ by $\tau(\mathcal{F})$. We say that a family $\mathcal{H}$ set-covers a family $\mathcal{F}$ if for each $F\in\mathcal{F}$ there exists $H\in\mathcal{H}$ such that $H\subset F$. For a graph $G$ denote $\mathcal{I}_{k}(G)$ the family of independent $k$-element sets of vertices of $G$. In the proofs of Theorems 1 and 4, we need the following result from [12]:
@@
We construct $\mathcal{H}$ inductively. Put $\mathcal{H}_{0}=\emptyset$. For each $l\in[0,\sqrt{k}-1]$, we construct a set-cover $\mathcal{H}_{l+1}\subset{[n]\choose l+1}$ for $\mathcal{F}$ from the set-cover $\mathcal{H}_{l}\subset{[n]\choose l}$ that we assume to have constructed. We also assume that $|\mathcal{H}_{l}|\leq k^{l}$. Put $\mathcal{F}[H]=\{F\in\mathcal{F}:H\subset F\}$. Using this notation, the family $\mathcal{H}_{l}$ being a set-cover means that $\mathcal{F}\subset\bigcup_{H\in\mathcal{H}_{l}}\mathcal{F}[H]$. Thus, it should be clear that $\mathcal{H}_{0}$ satisfies both the condition on size and on being a set-cover. For each $H\in\mathcal{H}_{l}$ consider a set $G_{H}\in\mathcal{G}^{\times}$ disjoint with $H$. Such set exists since $|H|=l<\sqrt{k}$ and $\tau(\mathcal{G}^{\times})\geq\sqrt{k}$. Put $\mathcal{H}_{l+1}=\{H\cup x:H\in\mathcal{H}_{l},x\in G_{H}\}$. Then it is easy to see that $|\mathcal{H}_{l+1}|\leq t|\mathcal{H}_{l}|\leq k^{l+1}$ and
$\mathcal{F}\subset\bigcup_{H\in\mathcal{H}_{l+1}}\mathcal{F}[H].$
@@ 2.2 Proof of Theorem 1
    $\mathcal{N}_{k}(\bar{G})\geq\frac{1+(2\gamma-1)(k-1)}{k}|V(\bar{G})|\mathcal{N}_{k-1}(\bar{G}),$
@@
    $\mathcal{N}_{3}(\bar{G})\geq\frac{1+(2\gamma-1)(3-1)}{3}|V(\bar{G})|\mathcal{N}_{2}(\bar{G}).$